MULTIPHOTON IONIZATION SPECTROSCOPY OF NO MOLECULE EMBEDDED IN LIQUID HELIUM DROPLET

\underline{ELENA POLYAKOVA}, DANIIL STOLYAROV and CURT WITTIG, Department of Chemistry, University of Southern California, Los Angeles, California 90089.

The Multiphoton Ionization spectrum of NO molecule trapped inside of the helium droplet was recorded for the first time in the UV-visible range.This spectrum shows severe broadening due the strong interaction of NO with helium superfluid environment. The Time of Flight spectra show the formation of NO^{+} (He_{N}) clusters with N = 0...40.